OpenStudy (nathaliama13):

A system of equations is shown below: n = 3m + 6 n − 2m = 2 What is the solution, in the form (m, n), to the system of equations? (−4, −6) (−5, −9) (2, 6) (1, 8)

OpenStudy (nathaliama13):

hi @paki do you think you can help me on this?

OpenStudy (nathaliama13):

@misty1212 can you help me?

OpenStudy (nathaliama13):

no one's helping me omg ;-;

OpenStudy (owlcoffee):

There are many ways to solve a system of equations, I'll show you the method most used when the system of equations have both variables in their composition, the method of "substitution". Let's call the equations, (1) and (2): \[1) n=3m+6\] \[2)n-2m=2\] Substitution, in a nutshell is isolating the variable in one equation and subtituting the variable on the other. But on equation (1) we can see that "n" is isolated on the left side of the equation, so that means that the variable is already solved for, so we can go ahead and replace it on equation (2): \[(3m+6)-2m=2\] And this is a first degree equation who has "m" as it's variable, and it's the only one, so we can solve for "m" and that should give us the value we are after, so we will operate similar terms: \[m+6=2\] \[m=2-6\] \[m=-4\] So, now that we have found the value of "m", you can take it and replace it in any of the two equations and then solve for "n" in order to obtain the value of "n" and completing the excercise.

OpenStudy (nathaliama13):

I solved it by elimination :-)